OpenStudy (anonymous):

How do I solve for x? 4x + 10yz = 0

OpenStudy (wolfe8):

You have 3 unknowns here. You can only solve for any unknown if you have 3 or more equations. What you can do with this though, is solve x in terms of the other unknowns. So do what you would do with any equality. Subtract both sides by 10yz: 4x + 10yz -10yz = 0 - 10yz 4x = -10yz Then divide both sides with 4: 4x/4= -10yz/4 x = -10yz/4 = -5/2yz = -2 1/2yz

OpenStudy (anonymous):

So not really finding x? Just finding x in terms of y and z?!

OpenStudy (wolfe8):

Yes. It's also called rewriting x in terms of y and z. You can only get a value for x if you have 3 equations as I said, or if you have y and z.